OpenStudy (anonymous):

Show that ∫_{0,∞}(dx/(1+x^n))=π/(nsin(π/n)) for ∀n∈ℝ. Repost, but I have no idea what to do, and the last one drowned.

OpenStudy (anonymous):

This?\[\int\limits \limits_{0}^{\infty} \frac{1}{1+x^{n}}dx = \frac{\pi}{n \sin\frac{\pi}{n}}\]

OpenStudy (anonymous):

Yep. This question was asked in context of the residue theorem http://en.wikipedia.org/wiki/Residue_theorem. Analysis and all that.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Residue_theorem The link, since the previous one apparently doesn't work. Not like it's too hard to Google it, or anything.

OpenStudy (asnaseer):

OK - I'll read up on the residue theorem and come back once I've understood it. Looks interesting.

OpenStudy (anonymous):

Thanks, asna!

OpenStudy (mr.math):

Is \(x\) a real variable?

OpenStudy (anonymous):

I don't believe I'm told if it is, but in context it's a safe assumption.

OpenStudy (turingtest):

is it? the wiki link suggest to me otherwise

OpenStudy (anonymous):

Point taken.

OpenStudy (mr.math):

I believe it's a real variable. Anyways, the relation above doesn't hold for \(n=1\) since \[\int_0^{\infty} \frac{1}{1+x}=\ln(1+x)|_0^{\infty} \to \infty.\]

OpenStudy (anonymous):

I'll be honest--I can't make heads or tails of the residue theorem.

OpenStudy (turingtest):

and I'm running out of tricks.... I think this is beyond me

OpenStudy (mr.math):

I'm not mastering the residue theorem, so I will first try to use it on the special case with \(n=2\), since we already know that the integral would be \(\frac{\pi}{2}\). \[\int_0^{\infty} \frac{1}{1+x^2}=\frac{1}{2}\int_{-\infty}^{\infty}\frac{dx}{1+x^2}=\frac{1}{2}\text{Re}\int_{-\infty}^{\infty}\frac{1}{1+z^2}dz.\] Now I will apply the theorem to evaluate \(\int_{-\infty}^{\infty}\frac{1}{1+z^2}dz\).

OpenStudy (asnaseer):

I found a you tube video where the person explains almost exactly this integral using the residue theorem: http://www.youtube.com/watch?v=XRFn0ioBPXI

OpenStudy (anonymous):

Whoa, nice find, asna. And clever algebra, math.

OpenStudy (asnaseer):

@badreferences: I tried to follow the reasoning in the video - it sort of makes sense but I will need to review some of his earlier videos in order to understand this completely. That will take me some time - so don't hold your breath! :) In the meantime, I'll leave you with the expert Mr.Math.

OpenStudy (anonymous):

I still don't see what you're trying to get at, though, Math. What can we determine by examining this special case? Looks like a fairly straightforward arctan.

OpenStudy (mr.math):

Yeah I know. I just wanted to do it using the residue theorem to see how it works with this simple case (that we already know its value). That's why I changed it to complex variable integral.

OpenStudy (mr.math):

This is a related video to the one that naseer posted above, where the man evaluated the integral of 1/(1+x^4) using the residue theorem. http://www.youtube.com/watch?v=MRLa5bk3_R4&feature=related

OpenStudy (mr.math):

I think it would help us understand the general case.

OpenStudy (anonymous):

Let's see if I can make myself understand this.